OpenStudy (anonymous):

The directions say simplify. I didn't know how to type this so over means like a fraction and the brackets should cover both upper and lower numbers. {-3(a^2b^-1)^3}^-2 over {7(a^-5b^6)^-1}

OpenStudy (anonymous):

Is this it? \[\frac{[-3(a^2b^{-1})^3]^{-2}}{7(a^{-5}b^6)^{-1}}\]

OpenStudy (anonymous):

No the brackets are over the bottom too so that the whole fraction is to the ^-2

OpenStudy (anonymous):

Ok so \[[\frac{-3(a^2b^{-1})^{3}}{7(a^{-5}b^6)^{-1}}]^{-2}\]

OpenStudy (anonymous):

Yes that's it

OpenStudy (anonymous):

So start with what's inside the brackets. Specifically work on the numerator. What do you have when you evaluate \[(a^2b^{-1})^3 = ?\]

OpenStudy (anonymous):

(a^6b^-3)

OpenStudy (anonymous):

Correct. Now the denominator.

OpenStudy (anonymous):

(a^-5b^-6

OpenStudy (anonymous):

Not quite.

OpenStudy (anonymous):

(a^-5b^6)

OpenStudy (anonymous):

Sorry I have to leave I'll have to come back to this later

OpenStudy (anonymous):

\[(a^{-5}b^6)^{-1} = a^{\text{-5 * -1}}b^{\text{6*-1}}\]